OpenStudy (anonymous):

If tan x° = and cos x° = , what is the value of sin x°? (1 points) sin x° = sin x° = sin x° = 10z sin x° = 10y

myininaya (myininaya):

\[\text{ if } \tan(x)=u \text{ and } \cos(x)=v \\ \text{ then recall } \tan(x)=\frac{\sin(x)}{\cos(x)} \\ \text{ so if you wanted \to solve for } \sin(x) \\ \text{ you multiply } \cos(x) \text{ on both sides } \sin(x)=\cos(x) \tan(x)\]

myininaya (myininaya):

so in this case sin(x)=u*v since sin(x)=cos(x)tan(x) where cos(x)=v and tan(x)=u